OpenStudy (hidden14):

I'm having trouble understanding exactly what an "amu" (or "u" or "Da" as I believe are more modern labels). What is a very layman explanation for "amu?"

OpenStudy (caozeyuan):

it is defined as 1/12 of a single carbon-12 atom, generally can be taken as the same as the mass of one proton or one neutron

OpenStudy (aaronq):

"amu" is an acronym for "atomic mass unit", 1 amu is defined as 1/12 of the mass of C-12. "Da" is a Dalton which means "g/mol", redundant but has been used historically for proteins

OpenStudy (caozeyuan):

Da, as in the abbreviation of Dalton, is the same unit as amu, but more common in the biochem lingo

OpenStudy (caozeyuan):

amu is the same as gram per mole is because number of atoms in 12 gram of C-12 is defined as one mole

OpenStudy (hidden14):

So then how would grams be converted to moles to atoms, or some combination of the two? Is it similar to some three variable equation like d=rt?

OpenStudy (aaronq):

\(\sf moles=\dfrac{mass}{Molar~mass}=\dfrac{\#~of~particles}{avogadro's~number }\) here "particles" can be a number of different things (atoms, molecules) depending on the context.

OpenStudy (hidden14):

@aaronq So for example, if I'm told to find the number of moles and atoms in 0.400g of Sodium, it would be\[\sf moles=\dfrac{mass}{Molar~mass}=\dfrac{0.400g~Na}{6.022x10^{23}}=6.64x10^{-25}\] and for atoms it would be: \[\sf atoms = moles * molar~mass = 6.64x10^{-25} * 6.022x10^{23} = 0.4~atoms\] Is my process here correct?

OpenStudy (aaronq):

You've made a mistake. You wrote 6.022*10^23 as the molar mass of Na which is incorrect.

OpenStudy (hidden14):

@aaronq I thought Avogadro's number was supposed to be there as it's in the denominator of the last quotient in your equation. Is this correct instead? \[\sf moles = \dfrac{mass}{Molar~mass}= \dfrac{0.400g Na}{22.99~amu} = 1.74*10^{-2}~moles\] \[\sf atoms = moles * molar~mass = (1.74*10^{-2})*(6.022*10^{23})= 1.05*10^{22} ~atoms\]

OpenStudy (aaronq):

Its the denominator when particles are involved - not mass, however. What you wrote now is correct.

OpenStudy (hidden14):

@aaronq So if I am given a substance with a mass in grams (kilograms or any mass related SI unit), and I need the moles of that substance, I convert it to grams, and divide it my it's atomic mass unit, and if I am given anything that is not grams or unable to be represented in grams (atoms, molecules, apples, trees), then I would use particles over Avogadro number?

OpenStudy (hidden14):

And by "anything unable to be represented in grams" I'm thinking absolutely anything that can't be modeled as "200g Na" or more generically "(mass in grams) of (substance)." Examples being: atoms, molecules, headphones, calculators, iPhones

OpenStudy (aaronq):

yup, that is correct.